An object moves along a straight line so that after 𝑑 minutes, its distance from its starting point

is 𝑠(𝑑) = 10t + (5/t+1) - 5 meters.

(a) At what speed is the object moving at the end of 4 minutes?

(b) How far does the object actually travel during the 5th minute?

(c) How fast is its velocity changing at 𝑑 = 6 π‘šπ‘–π‘›?


1
Expert's answer
2020-12-14T10:09:45-0500

"S(t)=10t+\frac{5}{t+1}-5\\
V(t)=S'(t)=10-\frac{5}{(t+1)^2}\\
\text{(a)}\\
V(4)=10-\frac{5}{(4+1)^2}\\
V(4)=10-\frac{5}{(5)^2}\\
V(4)=10-0.2\\
V(4)=9.8m/min.\\
\text{(b)}\\
S(5)-S(4)\\
(10(5)+\frac{5}{(5)+1}-5)-(10(4)+\frac{5}{(4)+1}-5)\\
=45.83-36\\
=9.83m.\\
\text{(c)}\\
V(t)=10-\frac{5}{(t+1)^2}\\
V(0)=10-\frac{5}{(0+1)^2}\\
=10-5\\
V(0)=5m/min\\
V(6)=10-\frac{5}{(6+1)^2}\\
V(6)=10-\frac{5}{(7)^2}\\
V(6)=10-0.10\\
V(6)=9.9m/min."